OpenStudy (anonymous):

A horizontal safely supports a load P that varies jointly as the product of the width W of the beam and the square of the depth D and inversely as the length L. A beam that has a width of 6 in, depth of 12 in, and length of 4 in can safely support 2,592 lb. Determine the safe load in pounds of a beam made from the same material if the beam is 12 in long. The safe load is?

OpenStudy (ribhu):

@ChloeSmith09 P = K. W.D^2 / L

OpenStudy (ribhu):

where K is constant of proportionality. and W,D and L are defined in question.

OpenStudy (ribhu):

2592 = K. 6.12.12/4 K=12

OpenStudy (ribhu):

now new P= 12.6.12.12/12 Pnew = 884 lbs

OpenStudy (ribhu):

this would be the answer